OpenStudy (anonymous):

Write the equation in slope-intercept form for 4x-10y=11

OpenStudy (anonymous):

still here geri?

OpenStudy (anonymous):

First thing to know is that the slope-intercept form looks like this: \[y=mx+b\].

OpenStudy (anonymous):

It is called the slope-intercept form because once you transform your equation into this form, m will be the slope and b will be the y-intercept of your line equation.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now all you have to do is solve for y in your given equation.

OpenStudy (anonymous):

Write out your work here and I'll be back to check on ya

OpenStudy (anonymous):

I got y=2/5x-11

OpenStudy (anonymous):

not bad, but a minor error. When you divided both sides by 10, the final step in solving for y, you should have divided it by ALL the terms on the other side. Think of them as being in parentheses if that helps.

OpenStudy (anonymous):

Do you follow?

OpenStudy (anonymous):

nope, is this answer wrong?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You moved the 10 from the y by dividing both sides of the equation by 10, so you should divide ALL the terms on both sides of the equation by 10. That includes 11. Do you see where you went wrong?

OpenStudy (anonymous):

y=2/5x+11

OpenStudy (anonymous):

Got it! y=2/5x -11/10...i think :(

OpenStudy (anonymous):

hey hey, some some confidence in yourself! you got it! Good job.

OpenStudy (anonymous):

phew! Thank you! God Bless your patience!